OpenStudy (anonymous):

Show that the function F of x equals the integral from 2 times x to 5 times x of 1 over t dt is constant on the interval (0, +∞).

OpenStudy (anonymous):

https://gyazo.com/e5425cec35c6e331d016d9664296b104

OpenStudy (anonymous):

@terenzreignz @zepdrix I used the fundamental theorem of calc to find derivative which was equal to 0.

OpenStudy (anonymous):

That's enough to prove it right?

terenzreignz (terenzreignz):

Use the second Fundamental Theorem instead.

OpenStudy (anonymous):

Yes That's what was done :p

OpenStudy (anonymous):

2nd, 1st it's all the same it's not a race ;)

OpenStudy (anonymous):

Slope=0 means no change in function

OpenStudy (owlcoffee):

It is not the same, the second part of the fundamental theorem for integral calculus is known also as Barrows rule, models the following equality: \[\int\limits_{a}^{b}f(x)dx=F(b)-F(a)\]

terenzreignz (terenzreignz):

I suppose I could start over, like I usually do :/ \[\Large \int\limits_{2x}^{5x}\frac1t dt \]

terenzreignz (terenzreignz):

And yes, did you use the FTC as Owlcoffee stated it above ^ ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I used

terenzreignz (terenzreignz):

Show me >:)

OpenStudy (anonymous):

to be exact

OpenStudy (anonymous):

\[F'(x)=u'(x)f(u(x))-v'(x)f(v(x))\]

OpenStudy (anonymous):

\[\int\limits_{v(x)}^{u(x)}f(t)dt\]

terenzreignz (terenzreignz):

It seems I may have to explain the second FTC after all :)

OpenStudy (anonymous):

O_o why?

OpenStudy (anonymous):

My teacher gave me that

OpenStudy (anonymous):

Is it the way I wrote it? Cause I was too lazy to write out the complete thing

OpenStudy (anonymous):

It has d/dx in front of the integral since that's what you need to do in this case

terenzreignz (terenzreignz):

Well, *I* am giving this to you now: Suppose you have a function F(x) such that the derivative of F is f on a given interval. That is: F'(x) = f(x) Then \[\Large \int\limits_a^b f(x) = F(a) - F(b)\] Is the gist of it.

OpenStudy (anonymous):

Find the derivative.

OpenStudy (anonymous):

Ok

terenzreignz (terenzreignz):

Things don't seem to be going my way ^^ \[\Large \int\limits_a^b f(x)dx = F(b) - F(a)\]

terenzreignz (terenzreignz):

There, use that. Now, can you integrate 1/t?

OpenStudy (anonymous):

ln(5x)-ln(2x)?

terenzreignz (terenzreignz):

That's right ^^ Simplify it using the rules concerning logarithms: I'd tell you which rule exactly, but maybe you already know :D

OpenStudy (anonymous):

A refresher would be nice

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

Just remembered

OpenStudy (anonymous):

\[\ln \frac{ 5x }{ 2x }\]?

terenzreignz (terenzreignz):

Yup. Simplify further?

OpenStudy (anonymous):

ln (5/2)

terenzreignz (terenzreignz):

Which is constant. Et voila ^^

OpenStudy (anonymous):

So is what I did considered wrong?

terenzreignz (terenzreignz):

I'm not exactly sure what you did :> But you do have that tendency to overcomplicate things sometimes :D But that which you used looked nothing like what was needed in this particular question... unless I'm mistaken.

OpenStudy (anonymous):

My teacher complicates it not me :p I'm using her formula.

terenzreignz (terenzreignz):

Well, that's why I'm here :D

OpenStudy (anonymous):

https://gyazo.com/cb41c86a344c6fb6dbadfadabd886845

OpenStudy (anonymous):

That's what I was given.

OpenStudy (anonymous):

I used a variant of the 1st which may seem complicated but it was plug and play as well.

terenzreignz (terenzreignz):

Part II comes to mind...

OpenStudy (anonymous):

It seems simpler but I don't feel like erasing and rewriting

OpenStudy (anonymous):

So I think I'm good with what I wrote

OpenStudy (anonymous):

:)